OpenStudy (luckylu):

How do I solve to find the x and y intercepts? y = x2 + x − 6? x-intercepts: (3, 0) and (2, 0) y-intercept: (0, 6) x-intercepts: (3, 0) and (−2, 0) y-intercept: (0, −6) x-intercepts: (−3, 0) and (2, 0) y-intercept: (0, −6) x-intercepts: (−3, 0) and (2, 0) y-intercept: (0, 6)

OpenStudy (3mar):

May I help?

OpenStudy (luckylu):

yes plz

OpenStudy (3mar):

Thanks a lot.

OpenStudy (3mar):

Do you want to know that algebraically or graphically?

OpenStudy (luckylu):

it's graph

OpenStudy (3mar):

As you wish. Could you plot this function and show me how does it look like?

OpenStudy (luckylu):

This is the entire question hold on.What are the x- and y-intercepts of the graph of y = x^2 + x − 6? x-intercepts: (3, 0) and (2, 0) y-intercept: (0, 6) x-intercepts: (3, 0) and (−2, 0) y-intercept: (0, −6) x-intercepts: (−3, 0) and (2, 0) y-intercept: (0, −6) x-intercepts: (−3, 0) and (2, 0) y-intercept: (0, 6) Litterally there is nothing else.

OpenStudy (3mar):

I know, Ms! If you don't mind I need you to graph such equation on the grid! 

OpenStudy (luckylu):

oh my bad

OpenStudy (3mar):

No problem, sister. Can you plot it?

OpenStudy (luckylu):

Idk how...

OpenStudy (3mar):

https://www.desmos.com/calculator/akcrnqgl4e like that for example!

OpenStudy (luckylu):

No and yes in a way it wants me to find what the intercepts are from the equation y= x^2+x-6

OpenStudy (3mar):

For y-intercept: it is logic that x is equal to zero, then put x=0 and tell what y is!

OpenStudy (luckylu):

-6 I got that

OpenStudy (luckylu):

idk how to get x

OpenStudy (luckylu):

It's one of the two middle choice answers B or C idk which. Since I can't figure out how to get x when there are two x's

OpenStudy (3mar):

-6 I got that \[\Huge\color{MediumBlue }\checkmark\] Yes, that is correct for Y. The same udea is for X. Let y=0 and let me know what you got!

OpenStudy (luckylu):

I could only get this and i don't think it's right... \[6=x^2+x\]

OpenStudy (3mar):

can you take x as a common factor from the right-hand side?

OpenStudy (luckylu):

would it become 1 or 2?

OpenStudy (3mar):

Can you share your works/steps, please?

OpenStudy (luckylu):

6= x^2+x 6= x^2+x -1 -1 5= x^2?

OpenStudy (jiteshmeghwal9):

To get the x intercept put y=0 To get the y intercept put x=0

OpenStudy (3mar):

To get x-values: form a full quadratic equation.